The client of a project has requested the project team to crash 8 hours of time....

  1. The client of a project has requested the project team to crash 8 hours of time. The table below provides the necessary information about the possible project crashing.

(a) What is the crash cost for 8 hours of time-savings?

(b) Assume the team calls the client and asks for a project extension, reducing the amount of time they need to crash. The project team has a $20 crash budget. Is the budget sufficient to crash 4 hours of time?

Activity

Normal

Duration (hours)

Normal

Cost ($)

Crash

Duration (hours)

Crash

Cost (S)

Immediate

Predecessors

A

2

10

2

0

None

B

3

15

2

23

A

C

5

25

4

30

B

D

3

20

1

24

C

E

6

30

4

45

C

F

1

5

1

0

E

G

7

35

6

50

F

H

10

50

7

80

D, G
Answer #1
Activity Crash cost per hour Allowable crashing hours
NA 0
B =(23-15)/(3-2) = 8 1
C =(30-25)/(5-4) = 5 1
D =(24-20)/(3-1) = 2 2
E = (45-30)/(6-4) = 7.5 2
F NA 0
G =(50-35)/(7-6) = 15 1
H =(80-50)/(10-7)= 10 3

Crash cost per hour = (Crash cost-Normal cost)/(Normal time- crash time)

Crash cost per hour tells us which activity is cheap to crash. The lower crash cost per hour the better.

If we have to crash project for 8 hours, we'll crash in Following order:

1.D by 2 hours (Cost =2*2 = 4)

2. C by 1 hour (Cost =1*5 = 5)

3. E by 2 hours (Cost =2* 7.5= 15)

4. B by 1 hour (Cost =1*8=8)

5. H by 2 hours (Cost = 2*10= 20)

Total crashing cost= 4+5+15+8+20 = 42

b.

For crashing by 4 hours, we have to crash-

D by 2 hours Costing 4, C by 1 costing 5 & E by 1 hour costing 7.5.

Total cost = 4+5+7.5 = 16.5

Total budget = 20

Therefore, budget is sufficient to crash project by 4 weeks.